OpenStudy (anonymous):

sqrt((13^2-(3^2+4^2))/(2*18)) The answer is 2, but how do you get it? Thank you

OpenStudy (anonymous):

\[\sqrt((13^2-(3^2+4^2))/(2*18))\]

OpenStudy (anonymous):

The whole thing is under the radical, correct? And if so, then 13^2-25 is 144. Aslo, 18*2=36. So what is under your radical is \[\sqrt{\frac{ 144 }{ 36 }}=\frac{ 12 }{ 6 }=2\]

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador